PREPARING ARTICLES WITH L^{A}TEX

INSTRUCTIONS TO AUTHORS FOR PREPARING COMPUSCRIPTS

ELSEVIER

SCIENCE

PUBLISHERS B.V.

PREPARING ARTICLES WITH L^{A}TEX

INSTRUCTIONS TO AUTHORS FOR PREPARING COMPUSCRIPTS

ELSEVIER SCIENCE PUBLISHERS B.V.

This publication was typeset using L^{A}TEX

TEX is a trademark of the American Mathematical Society

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994 by Elsevier Science Publishers B.V. All rights reserved Contents

1 Introduction 5 2 Preparing a compuscript 6 2.1 Title and author 6 2.2 Simple text 7 2.3 Sectional units 8 2.4 Lists 8 2.5 Cross-references 8 2.6 Mathematical formulas 9 2.7 Theorems and definitions 11 2.8 Proofs 12 2.9 Literature references 12 2.10 Tables and figures 13 2.11 Programs and algorithms 13 2.12 Large articles 14 2.13 Private definitions 14 2.14 Layout 14 2.15 Deviations from standard document styles 15 2.16 Technical information, and versions of L^{A}TEX 15 3 Submitting a compuscript 17 3.1 Sending via electronic mail 17 3.2 Submission on diskette 18 4 Getting help 18 A Examples 20 4 1. Introduction

Nowadays, it is becoming more and more customary for authors to type their manuscripts using some kind of electronic device and composing the result with some text-processing system. Systems that are quite popular are TEX and L^{A}TEX. In order to assist authors in preparing their papers for articles published by Elsevier Science Publishers in such a way that their files can be used to print the article from, we have developed L^{A}TEX document styles for our journals. The following is a description of these document styles. For best understanding, authors should be reasonably familiar with the L^{A}TEX manual written by Leslie Lamport [1].

In order to enable the publisher to bring the article into the uniform layout and style of the journal in which it will appear, authors are kindly requested to follow the suggestions mentioned below. This has the advantage of keeping editorial changes to a minimum, which will considerably speed up the publication process.

Upon receipt of the compuscript, it is given to a technical editor, who prints the compuscript on paper, reads it carefully and makes changes when necessary. If sending proofs is part of the normal procedure for the particular journal, a proof is sent to the author. If the author finds something in the proof that should be changed, he/she should indicate this clearly in the margin, so that the technical editor can apply these corrections before making the paper ready for publication.

For all journals that accept author-prepared L^{A}TEX articles we have document styles. All these document styles, which are used for the actual production of the journals, have the same commands. Furthermore, there is a separate document style elsart that is fully compatible with the production document styles. Authors can use this document style elsart to obtain preprint output. When the article is prepared for publication, this document style is replaced by a document style for the journal in which the article is published.

This documentation contains a user’s guide, guidelines for submitting the article for publication and information on where to get help in case problems occur.

2. Preparing a compuscript

The documentstyle elsart, with which the article can be prepared and preprint output can be obtained, is compatible with the standard document styles of L^{A}TEX, except for the specification of the front matter, i.e. the title, author, addresses and abstract.

In the following sections we will describe the differences between normal L^{A}TEX usage and the usage of the Elsevier document styles. Also, we will summarize some of the important aspects of coding a compuscript with L^{A}TEX.

2.1. Title and author

In the Elsevier document styles the commands \title, \author etc., have been replaced by a more general frontmatter environment. Since the standard L^{A}TEX document styles do not differentiate between author name and address, extra mark-up instructions have been added to the Elsevier document styles. Within the frontmatter environment, you should specify the title, names and addresses of the authors, followed by the abstract and – in some cases – a keyword abstract.^{1} Title, author, collaboration, address, abstract and keyword abstract should be indicated with the instructions

\title, \author, \collab and \address, and the abstract and keyword

environments, respectively. The instruction \maketitle has become obsolete in the Elsevier document style.

There are two types of author–address lists. These are illustrated by Examples 1 and 2 in Appendix A. The first type of author–address list consists of one or more groups of authors followed by an address (affiliation). In this type of list there is an implicit link between authors and addresses. The second type of author–address list consists of one list of all authors, followed by one list of all addresses (affiliations), and with explicit links between authors and addresses. The links are written as optional arguments to the \author, \collab and \address commands and are usually formatted as footnote-like symbols.

The \thanks command can be used to produce notes that are added to the title, author or address. In the Elsevier document styles this command should be written inside the frontmatter environment, but outside the argument of \title, \author, \collab and \address; see also Examples 1 and 2. The modified \thanks command has an optional argument that can be used to attach a label to a note:

\thanks[CAICYT]{Partially supported by CAICYT, Spain.}

1 Optional, not present in some journals.

Inside the argument of \title, \author, \collab and \address one can refer to this note with the command \thanksref, which takes the label of a \thanks command as argument:

\author{L.A. Fernandez\thanksref{CAICYT}}

The command \and has its usual meaning.

In some journals, authors of experimental papers have to add keyword abstracts. These abstracts are specified by using an equivalent of the abstract environment: the keyword environment. The following input gives an example of the use of this environment.

\begin{keyword}
Radioactivity.
($\beta^+$, EC) [from Pt(p, $x$n)Au or ...
\end{keyword}

might generate this output

Keywords: Radioactivity. (β^{+}, EC) [from Pt(p, xn)Au or ...

The proper position of the keyword environment is inside the frontmatter environment, before or after the abstract environment.

2.2. Simple text

Text should be typed as usual. Hyphens are typed as -, number ranges are typed as --. The en dash -- is also used, e.g., in ‘Theorem of Cantor– Schröder–Bernstein’.

Emphasized text is obtained with the command \em. In most cases this will result in italic text representing emphasis. Italic text should be terminated by an italic correction, i.e.

{\em heavy quarks\/}

unless the text in italics is immediately followed by a full stop (.) or comma (,).

Extra or exceptional hyphenations are added to TEX’s list of abbreviated words by means of the command \hyphenation, which should be placed in the preamble of the document. An example:

\hyphenation{caus-al min-i-mi-za-tion pro-ven}

Introduce macros (with care, see 2.13) for notations and abbreviations that occur more than once, for example ‘e.g.’ and ‘i.e.’. This facilitates changes in notation. If you introduce macros for abbreviations, these are often parameterless macros, so you should be aware of TEX’s behaviour with regard to spaces following a parameterless macro. An instruction without parameters should be defined and used as

\newcommand{\ie}{i.e.}
...
... extra particles,
\ie{} particles ...

... extra particles, i.e. particles ...

Alternatives to \ie{} are \ie\␣ and {\ie}. The \␣ after \ie produces a space, whereas \ie␣particles will result in ‘i.e.particles’ [1, p. 16].

Putting a space in the definition of \ie is not the right solution, since it can result in a space before a punctuation mark, e.g.

\newcommand{\ie}{i.e. }
...
... extra particles,
\ie, particles ...

... extra particles, i.e. , particles ...

2.3. Sectional units

Sectional units are obtained in the usual way, i.e. with the L^{A}TEX in-

structions \section, \subsection, \subsubsection, \paragraph and
\subparagraph.

A new environment ack – see also Section 2.7 – has been added to produce an ‘Acknowledgements’ section, which should be placed at the end of the article, just before the references.

2.4. Lists

Lists of items are produced with the usual itemize and enumerate environments. The itemize environment is used for unnumbered lists and the enumerate environment for numbered lists. Even if the layout of these lists is not precisely what you would like, we prefer lists to be coded this way instead of by hand. This enables the document style for the specific journal to determine the list layout.

2.5. Cross-references

Use \label and \ref for cross-references to equations, figures, tables, sections, subsections, etc., instead of plain numbers. For references to the literature list at the end of the article see Section 2.9.

Every numbered part to which one wants to refer, should be labelled with the instruction \label. For example:

\begin{equation}
  e^{i\pi} + 1 = 0   \label{eq:euler}

\end{equation}

With the instruction \ref one can refer to a numbered part that has been labelled:

..., see also eq. (\ref{eq:euler})

The \label instruction should be typed

• immediately after (or one line below), but not inside the argument of a number-generating instruction such as \section or \caption, e.g.:

     \caption{Cross section}  \label{fig:crosssec}

• roughly in the position where the number appears, in environments such as equation, e.g.:

     \begin{equation}
       e^{i\pi} + 1 = 0   \label{eq:euler}
     \end{equation}

2.6. Mathematical formulas

For in-line formulas use \( ... \) or $ ... $. Avoid built-up constructions, for example fractions and matrices, in in-line formulas.

For unnumbered displayed one-line formulas use the displaymath environment or the shorthand notation \[ ... \]. For numbered displayed oneline formulas use the equation environment. Do not use $$ ... $$, but only the L^{A}TEX environments, so that the document style determines the formula layout. For example, the input for:

(P + \frac{a}{V^{2}} ) (V − b) = RT,  (1)

is:

\begin{equation}
\left( P + \frac{a}{V^2} \right) (V-b) = RT ,
\end{equation}

For displayed multi-line formulas use the eqnarray environment. For example,

\begin{eqnarray}
f(x) & = & \sum_{n=1}^{\infty} a_n\cos(nx) +
                               b_n\sin(nx) \nonumber \\
     & = & \sum_{n=-\infty}^{\infty}
                            c_n\exp(-\mathrm{i} xn)\, .
\end{eqnarray}

produces:

f(x) = ∑_{n=1}^{∞} a_{n} cos(nx) + b_{n} sin(nx)

= ∑_{n=−∞}^{∞} c_{n} exp(−ixn) .  (2)

Angle brackets, which are used in, e.g., the inner product notation, the ‘braket’ notation (physics), and in BNF (computer science), are obtained with

\langle and \rangle:
\langle x, y   \rangle = 0
\langle p|A|p' \rangle = 0
\langle \mbox{sign} \rangle
\longrightarrow + | -

⟨x, y⟩ = 0 ⟨p|A|p′⟩ = 0 ⟨sign⟩ ⟶ +|−

Superscripts and subscripts that are words or abbreviations, as in σ_{low}, should be typed as roman letters; this is done as follows:

\( \sigma_{\mathrm{low}} \)  σ_{low}

instead of

\( \sigma_{low} \)  σ_{low}

The most common symbols that are conventionally typeset in a roman typeface, for example units, are listed below. For some of these, see also Table 1 on page 16.

• The Euler number, for example, e^{x}.

• i when used as imaginary unit, e.g. a+bi or e^{iϕ}, etc. The Euler equation, which was used as an example earlier, can therefore also be typed as

     \begin{equation}
       \mathrm{e}^{\mathrm{i}\pi} + 1 = 0
                                \label{eq:euler}
     \end{equation}

• Geometric functions, e.g. exp, sin, cos, tan, etc. L^{A}TEX provides macros \sin, \cos, \tan for these and similar functions. These macros also give the proper spacing in mathematical formulas.

• The differential operators, e.g. dx, and the operators Im and Re for the imaginary and real parts of complex numbers, respectively.^{2}

• Groups, for example SU(2) and SU(3).

• Labels for atomic orbitals and atomic shells. Example: 4s, 4p, K, L.

2 The normal shape of Greek capital letters is upright. The slanted shape of, e.g., the letter Δ is obtained with \varDelta, as in AMS-L^{A}TEX: Δ.

• Greek letters when used as a unit, e.g. Ω for ohm.

• Units in general. Example: cm, Å, and b for barn.

• Subscripts and superscripts that are used as an abbreviation. Examples: T_{C} (Curie temperature), T_{c} (critical temperature), and C_{3v} (identifier of space group)

• Operator or function names, or their abbreviatons, e.g. Ker, Im, Hom, Re, etc.

Of the advanced features of TEX we mention the possibility to define extra symbols. Extra relation symbols can be defined as in the following example (see also Section 2.13):

\newcommand{\leL}{\mathrel{\le_{\mathrm{L}}}}
\( a \leL b \)

produces the following result:

a ≤_{L} b

Extra log-like functions or operators can be defined as follows:

\newcommand{\re}{\mathop{\mathrm{Re}}}
\newcommand{\im}{\mathop{\mathrm{Im}}}
\( z + \bar{z} = 2 \re z,  \quad
   z - \bar{z} = 2 \mathrm {i} \im z \)

produces the following result:

z + z̄ = 2 Re z, z − z̄ = 2i Im z

For more information on TEX’s advanced mathematical features we refer to chapters 16–18 of the TEX book [3]. It is also possible to use the AMS-L^{A}TEX package [4], which can be obtained from the AMS, from various TEX archives, or from us (see Section 4).

2.7. Theorems and definitions

L^{A}TEX provides \newtheorem to create theorem environments. The Elsevier document styles contain a set of pre-defined environments for theorems, definitions, proofs, remarks and the like.

The following environments are defined (analogous to the example given in the AMS-L^{A}TEX user’s guide [4, §31.5]):

Environment name Heading  Environment name Heading
thm  Theorem  exmp  Example
lem  Lemma  prob  Problem
cor  Corollary  rem  Remark
prop  Proposition note  Note
crit  Criterion  claim  Claim
alg  Algorithm summ  Summary
defn  Definition case  Case
conj  Conjecture ack  Acknowledgement

To add theorem-type environments to an article, use the \newtheorem command – see the L^{A}TEX user manual [1].

2.8. Proofs

The Elsevier document styles also provide a predefined pf environment, and a starred form pf*, for proofs. The pf environment produces the heading ‘Proof’ with appropriate spacing and punctuation. A ‘Q.E.D.’ symbol, □, can be appended at the end of a proof with the command \qed.

The starred form, pf*, of the proof environment takes an argument in curly braces, which allows you to substitute a different name for the standard ‘Proof’. If you want to substitute, say, ‘Proof (sufficiency)’, then write

\begin{pf*}{Proof (sufficiency)}

2.9. Literature references

The list of literature references can be produced in two ways, by using

• the environment thebibliography, or

• BibTEX Example 3 shows a bibliography produced with the thebibliography environment.

If the references are collected in one, not too large, BibTEX file (.bib), it would be appreciated if you would let us have this file as well. In a future release we will include a BibTEX bibliography style in the author package as well.

The instruction \cite should be used to obtain references to this list, i.e. citations. The Elsevier document styles take care of the actual formatting of the citation, e.g. as roman numbers between brackets, or as a superscript number.

For multiple citations do not use \cite{Knuth}\cite{Lamport}, but use \cite{Knuth,Lamport} instead. Consecutive numbers in a citation appear as a range, i.e. [1,2,3] is automatically converted by the document style to [1–3]. For a note added to a citation use \cite[note]{key}, for example:

\cite[p. 217]{Knuth}.

2.10. Tables and figures

Put the tables and figures in the text with the table and figure environments, and position them near the first reference of the table or figure in the text. Do not put them at the end of the article.

A figure is obtained with

\begin{figure}
\vspace{30mm} % height of figure
\caption{ ... text below figure ... }
\end{figure}

Instead of the instruction \vspace{30mm} for the white space to be reserved for a separate figure, you can insert diagrams. Simple diagrams can be drawn with the picture environment; Feynman diagrams with the feynman package. Originals of separate figures should be sent via ordinary mail; sometimes good-quality prints are also acceptable. These should be submitted in the usual way. We can also process figures in PostScript form, but not in any other electronic form.

A table is obtained with

\begin{table}
\caption{ ... text above table ... }
\begin{tabular}{ ... }
...
\end{tabular}
\end{table}

Please avoid long caption texts – in figures and tables – and vertical rules.

2.11. Programs and algorithms

Fragments of computer programs and descriptions of algorithms should be prepared as if they were normal text. Use the same fonts for keywords, variables, etc., as in the text; do not use small typeface sizes to make program fragments and algorithms fit within the margins set by the document style.

An example, with only the tabbing environment and one new definition:

\newcommand{\keyw}[1]{{\bf #1}}
\begin{tabbing}
\quad \=\quad \=\quad            \kill
\keyw{for} each $x$ \keyw{do}    \\
\> \keyw{if} extension$(p, x)$   \\

\> \> \keyw{then} $E:=E\cup\{x\}$\\
\keyw{return} $E$
\end{tabbing}

This produces the following:

for each x do

if extension(p, x)

then E := E ∪ {x} return E

2.12. Large articles

A compuscript can be submitted as one or more files. If there is more than one file, one of them should be a root file. The root file inputs the files that constitute the entire article by means of \input or \include.

2.13. Private definitions

Private definitions should be placed in the preamble of the article, and not at any other place in the document. Such private definitions, i.e. definitions

made using the commands \newcommand, \renewcommand, \newenvironment

or \renewenvironment, should be used with great care.

Sensible, restricted usage of private definitions is encouraged. Large macro packages should be avoided. Definitions that are not used in the article should be omitted. Do not change existing environments, commands and other standard parts of L^{A}TEX. Definitions that are merely abbreviations for keystrokes, such as \bt for \begin{theorem}, should be avoided (use the facilities of your editor program to minimize keystrokes). A short description of the various definitions, in the form of TEX comment lines, is appreciated.

Deviation from these rules may cause inaccuracies in the article or a delay in publication, or may even result in the L^{A}TEX file being discarded altogether so that the article is typeset conventionally.

2.14. Layout

The document style elsart, which is part of this package, can be used to obtain preprint output. When the article is prepared for publication, this document style is replaced by a document style for the journal in which the article will be published.

The elsart style is compatible with all Elsevier’s journal styles, so that preparation of the article for final publication is straightforward.

In order to facilitate our processing of your article, please give easily identifiable structure to the various parts of the text by making use of the usual L^{A}TEX commands or by your own commands defined in the preamble, rather than by using explicit layout commands, such as \hspace, \vspace, \large, \centering, etc. Also, do not redefine the page-layout parameters.

2.15. Deviations from standard document styles

The document style elsart deviates from the standard document styles in the following areas

• specification of the front matter

• extra commands The document style defines several extra instructions. These are summarized in Table 1.

The document style redefines the standard command \vec: it formats vector symbols according to the layout of the journal, often italic boldface letters. The command \pol produces the standard vector notation, i.e. with a small right arrow on top of the argument.

2.16. Technical information, and versions of L^{A}TEX

In June 1994 a new version of L^{A}TEX was released, L^{A}TEX 2ε; Elsevier will continue to support users of the old L^{A}TEX209 for the foreseeable future, but would like authors to switch to L^{A}TEX 2ε as soon as practical. It is documented in the second edition of Lamport’s book [?], and described in great detail in [2].

Our preprint style is available in two forms, as eslart.sty and eslart.cls. The document style elsart.sty, with the corresponding elsart12.sty has been designed for L^{A}TEX 2.09 (version of January 1992 or later). The document class eslart.cls (no extra size file) has been designed for L^{A}TEX2e (versions from June 1994).

It is also possible to use the document style or class in combination with the AMS-L^{A}TEX package [4], in its L^{A}TEX209 or L^{A}TEX 2ε version, and we recommend this to authors who have more complex mathematical needs.

Table 1: Extra commands.

Front matter commands

\title{string}  title of article
\author[key]{string} name of one author
\collab[key]{string} name of collaboration (group of authors)
\address[key]{string} address of author or collaboration
\thanks[key]{string} note to one of the above elements
\thanksref{key}  reference to \thanks note

Case fractions

\half  small \frac{1}{2}

\threehalf  small \frac{3}{2}

\quart  small \frac{1}{4}

Theorem environments

–  see Sections 2.7 and 2.8

Extra mathematical operators

\d  differential ‘d’
\e  base of natural logarithm
other operators  see below
Blackboard bold symbols (AMSFonts version 2.1 must be present)
\Nset  ℕ, set of positive integer numbers
\Zset  ℤ, set of integer numbers
\Qset  ℚ, set of rational numbers
\Rset  ℝ, set of real numbers
\Cset  ℂ, set of complex numbers
\Hset  ℍ, set of quaternions
other letters  use \Bbb{...} from amsfonts

Extra notations for physics

\nuc  nuclides, \nuc{183}{Ir} produces ‘^{183}Ir’
\vec  boldface vector
\pol  polarization (right arrow on top of argument)
\FMslash  small slash through letter (Feynman notation)
\FMSlash  large slash through letter (Feynman notation)

3. Submitting a compuscript

The guidelines for submission of compuscripts can be found on the inside cover pages of the journal to which you would like to submit the article. If submission via electronic mail is allowed, you will find the network address to which you can send your compuscript in those guidelines as well.

For passing a compuscript to the publisher for final processing we give the following guidelines.

3.1. Sending via electronic mail

Short articles (say, less than 30 pages) should be prepared as one L^{A}TEX file and be sent via electronic mail as one message. Large files may be split into several parts, which are input in the root file.

• Send all files in separate messages; do not concatenate them together in one large message.

• Identify each part in the subject line as ‘part m of n’ in addition to the identification described above; note that without proper identification the order of the parts will be lost in the mail.

• If the article consists of more than five files, we prefer submission on diskette (see below) or via FTP. Please contact the publisher for more information on the latter.

If you send your compuscript via electronic mail, prepare the file such that no line is longer than 72 characters. This also prevents loss of information in various networks. Include

1. name of sender,

2. journal identification and article number, and

3. name of the file in the subject line of your electronic-mail message.

Also, include an ASCII table at the start of every file you send via electronic mail. An ASCII table, filename ascii.tab, is part of the package authors can obtain from the publisher. It contains the following:

% Upper-case    A B C D E F G H I J K L M N O P Q R S T U V W X Y Z
% Lower-case    a b c d e f g h i j k l m n o p q r s t u v w x y z
% Digits        0 1 2 3 4 5 6 7 8 9
% Exclamation   !           Double quote "          Hash (number) #
% Dollar        $           Percent      %          Ampersand     &
% Acute accent  '           Left paren   (          Right paren   )
% Asterisk      *           Plus         +          Comma         ,
% Minus         -           Point        .          Solidus       /
% Colon         :           Semicolon    ;          Less than     <
% Equals        =           Greater than >          Question mark ?

% At            @           Left bracket [          Backslash     \
% Right bracket ]           Circumflex   ^Underscore    _
% Grave accent  `           Left brace   {          Vertical bar  |
% Right brace   }           Tilde        ~

If this is included, any distortion can be detected and removed from the submitted files.

3.2. Submission on diskette

If you submit your compuscript on a diskette, prepare the file such that no line is longer than 72 characters. Try to use as few diskettes as possible, and put a label, with

1. name of sender, and

2. journal identification and article number

on each of them. Also add a file readme with a list of all the files on the diskettes and a description of their contents.

The allowed diskette types are: MS-DOS 3.5 inch, MS-DOS 5.25 inch and Macintosh, and for every diskette type all densities are possible.

4. Getting help

Although a lot of effort has been put in keeping the document style easy to use and in obtaining a concise description of the most common aspects of style, it is of course possible that authors encounter problems while using it. Also authors might have suggestions for additions. In those cases they should send their comments and suggestions to the address mentioned on the inside cover of the journal.

References

[1] Leslie Lamport: L^{A}TEX, A document preparation system, 2nd edition, Addison-Wesley (Reading, Massachusetts, 1994)

[2] Michel Goossens, Frank Mittelbach and Alexander Samarin: The L^{A}TEXCompanion, Addison-Wesley (Reading, Massachusetts, 1994)

[3] Donald E. Knuth: The TEX book Addison-Wesley (Reading, Massachusetts, 1986)

[4] AMS-L^{A}TEX Version 1.1—User’s Guide, American Mathematical Society, Providence, R.I., December 1990; distributed with the AMS-L^{A}TEX package.

[5] Frank Mittelbach and Rainer Schöpf: The new font family selection— user interface to standard L^{A}TEX TUGboat 11 (1990) 297–305.

A. Examples

In this appendix we will show a few examples of the use of the document style elsart: two examples of the front matter, and one example of the bibliography environment. L^{A}TEX 2ε users should simply substitute

\documentclass in place of \documentstyle.

\documentstyle{elsart}
\begin{document}
\begin{frontmatter}
\title{Integrability in
       random matrix models\thanksref{talk}}
\thanks[talk]{Expanded version of a talk
  presented at the Singapore Meeting on
  Particle Physics (Singapore, August 1990).}
\author{L. Alvarez-Gaum\'{e}}
\address{Theory Division, CERN,
  CH-1211 Geneva 23, Switzerland}
\author{C. Gomez\thanksref{SNSF}},
\address{D\'{e}partment de Physique Th\'{e}orique,
  Universit\'{e} de Gen\`{e}ve,
  CH-1211 Geneva 4, Switzerland}
\author{J. Lacki},
\address{School of Natural Sciences,
  Institute for Advanced Study,
  Princeton, NJ 08540, USA}
\thanks[SNSF]{Supported by the
  Swiss National Science Foundation}
\begin{abstract}
We prove the equivalence between the recent matrix model
formulation of 2D gravity and lattice integrable models.
For even potentials this system is the Volterra hierarchy.
\end{abstract}
\end{frontmatter}
\section{Introduction}
Some aspects of the recently discovered non-perturbative
solutions to non-critical strings \cite{ref1} can be better
understood and clarified directly in terms of the
integrability properties of the random matrix model.
...

Example 1. Article opening with implicit links (input).

Integrability in random matrix models^{⋆}

L. Alvarez-Gaumé

Theory Division, CERN, CH-1211 Geneva 23, Switzerland

C. Gomez^{1}

Départment de Physique Théorique, Université de Genève, CH-1211 Geneva 4, Switzerland

J. Lacki

School of Natural Sciences, Institute for Advanced Study, Princeton, NJ 08540, USA

Abstract We prove the equivalence between the recent matrix model formulation of 2D gravity and lattice integrable models. For even potentials this system is the Volterra hierarchy.

1. Introduction

Some aspects of the recently discovered non-perturbative solutions to noncritical strings [1] can be better understood and clarified directly in terms of the integrability properties of the random matrix model.

...

⋆ Expanded version of a talk presented at the Singapore Meeting on Particle Physics (Singapore, August 1990). 1 Supported by the Swiss National Science Foundation

Example 1. Article opening with implicit links (output).

\documentstyle{elsart}
\begin{document}
\begin{frontmatter}
\title{A renormalization group study of a gauge \\
       theory: SU(3) at finite temperature}
\author[Madrid]{L.A. Fernandez\thanksref{CAICYT}},
\author[Pisa]{M.P. Lombardo},
\author[Rome]{R. Petronzio} and
\author[Zaragoza]{A. Tarancon\thanksref{CAICYT}}
\address[Madrid]{Departamento de F\'{\i}sica Te\'{o}rica,
  Universidad Complutense de Madrid, E-28040 Madrid, Spain}
\address[Pisa]{INFN, Sezione di Pisa, I-56100 Pisa, Italy}
\address[Rome]{Dipartimento di Fisica,
  Universit\`{a} di Roma II ``Tor Vergata'' and
  INFN, Sezione di Roma Tor Vergata,
  Via O. Raimondo, I-00173 Rome, Italy}
\address[Zaragoza]{Departamento de F\'{\i}sica Te\'{o}rica,
  Universidad de Zaragoza, E-50009 Zaragoza, Spain}
\thanks[CAICYT]{Partially supported by CAICYT, Spain.}
\begin{abstract}
We apply a finite size renormalization group method to the
study of the deconfining transition in pure gauge SU(3). By
constructing renormalized systems with $2^3$ and 2 variables
suitably defined we obtain a very accurate determination
of the transition point and of the thermal exponent $\nu$.
\end{abstract}
\end{frontmatter}
The pure gauge SU(3) system at finite temperature
undergoes a phase transition from the confined to
the deconfined phase associated to the spontaneous
breaking of the local Z(3) symmetry.
...

Example 2. Article opening with explicit links (input).

A renormalization group study of a gauge theory: SU(3) at finite temperature

L.A. Fernandez^{a,1}, M.P. Lombardo^{b}, R. Petronzio^{c}, A. Tarancon^{d,1}

a Departamento de Física Teórica, Universidad Complutense de Madrid,

E-28040 Madrid, Spain b INFN, Sezione di Pisa, I-56100 Pisa, Italy

c Dipartimento di Fisica, Università di Roma II “Tor Vergata” and INFN, Sezione di Roma Tor Vergata, Via O. Raimondo, I-00173 Rome, Italy d Departamento de Física Teórica, Universidad de Zaragoza, E-50009

Zaragoza, Spain

Abstract We apply a finite size renormalization group method to the study of the deconfining transition in pure gauge SU(3). By constructing renormalized systems with 2^{3} and 2 variables suitably defined we obtain a very accurate determination of the transition point and of the thermal exponent ν.

The pure gauge SU(3) system at finite temperature undergoes a phase transition from the confined to the deconfined phase associated to the spontaneous breaking of the local Z(3) symmetry.

...

1 Partially supported by CAICYT, Spain.

Example 2. Article opening with explicit links (output).

\begin{thebibliography}{9}
\bibitem{Robi66}
A. Robinson,
{\em Non-standard Analysis\/}
(North-Holland, Amsterdam, 1966).
\bibitem{Sand89a}
E. Sandewall,
Combining logic and differential equations
for describing real-world systems,
in: R.J. Brachmann, H. Levesque and R. Reiter, eds.,
{\em Proceedings First International Conference on
Principles of Knowledge Representation and Reasoning\/}
(Morgan Kaufmann, Los Altos, CA, 1989) 412--320.
\bibitem{Sand89b}
E. Sandewall,
Filter preferential treatment for the logic of action
in almost continuous worlds,
in: R.J. Brachmann, H. Levesque and R. Reiter, eds.,
{\em Proceedings IJCAI-89\/}
(Detroit, MI, 1989) 894--899.
\bibitem{Shoh88a}
Y.Shoham,
Chronological ignorance:
experiments in nonmonotonic temporal reasoning,
{\em Artif. Intell.\/} {\bf 36} (1988) 279--331.
\bibitem{Shoh88b}
Y.Shoham and D. McDermott,
Problems in formal temporal reasoning,
{\em Artif. Intell.\/} {\bf 36} (1988) 49--61.
\bibitem{Bent83}
J. van Benthem,
{\em The logic of time\/}
(Reidel, Dordrecht, 1983).
\end{thebibliography}

Example 3. Literature references (input).

References

[1] A. Robinson, Non-standard Analysis (North-Holland, Amsterdam, 1966).

[2] E. Sandewall, Combining logic and differential equations for describing real-world systems, in: R.J. Brachmann, H. Levesque and R. Reiter, eds., Proceedings First International Conference on Principles of Knowledge Representation and Reasoning (Morgan Kaufmann, Los Altos, CA, 1989) 412–320.

[3] E. Sandewall, Filter preferential treatment for the logic of action in almost continuous worlds, in: R.J. Brachmann, H. Levesque and R. Reiter, eds., Proceedings IJCAI-89 (Detroit, MI, 1989) 894–899.

[4] Y.Shoham, Chronological ignorance: experiments in nonmonotonic temporal reasoning, Artif. Intell. 36 (1988) 279–331.

[5] Y.Shoham and D. McDermott, Problems in formal temporal reasoning, Artif. Intell. 36 (1988) 49–61.

[6] J. van Benthem, The logic of time (Reidel, Dordrecht, 1983).

Example 3. Literature references (output).

Index

abbreviations

macros for, 7 abstract, 6, 7

abstract, 6

acknowledgements, 8 address, 6

\address, 6

optional argument of, 6 algorithm, 13

\and, 7

angle brackets, 10 author, 6

\author, 6

optional argument of, 6

bibliography made with BibTEX, 12

   made with bibliography en-

vironment, 12 BibTEX, 12

caption, 13 argument too long, 13 vertical rules, 13

\caption, 13

in combination with \label, 9 citation, 12

formatting the, 12

multiple, 13

with added note, 13

\cite, 12

collab, 6

\collab, 6

optional argument of, 6 computer program, 13 cross-reference, 8

dash, 7 diagram, 13 differences with standard styles, 6

emphasized text, 7

                                  enumerate, 8

eqnarray environment, 9 equation, 9

displayed, 9 in-line, 9 multi-line, 9 equation environment, 9 extra instructions, 15

figure environment, 13

                                      \label in, see \caption

formula, 9 displayed, 9 in-line, 9 multi-line, 9 front matter, 6, 7

hyphen, 7

                                  \include, 14
                                  \input, 14

italic correction, 7

italics, 7

                                  itemize, 8

keyword abstract, 6, 7

                                  \label, 8

for equation number, 9 for sectional unit, 9 for table or figure caption, 9

layout

explicit commands for, 15 lists, 8 literature references, 12

                                  \maketitle, 6

notations

macros for, 7 number ranges, 7

parameterless macro, 7

picture, 13

preamble, 7 proof environment, 12

\ref, 9

roman typeface, 10 root file, 14

\section

in combination with \label, 9 sectional units, 8 space

explicit, 8 submitting a compuscript

on a diskette, 18

via electronic mail, 17 subscripts

abbreviations in, 10

words in, 10 superscripts

abbreviations in, 10

words in, 10

table environment, 13

   \label in, see \caption
\thanks, 6

optional argument of, 6

\thanksref, 7

theorem environments, 11 title, 6

\title, 6

units, 10 user-defined

log-like functions, 11

operators, 11

relation symbols, 11

\vec, 15

vector, 15